# Machine Learning for Network Attacks Classification and Statistical Evaluation of Adversarial Learning Methodologies for Synthetic Data Generation

### 01 Apr 2026 v2

@@ II-A Data Collection
For all the results that will be presented in the next sections we used the under-sampled version of the UMNIDS dataset, a newly created dataset, which was created by processing raw pcap files from CIC-IDS 2017, CIC-IOT 2023, UNSW-NB15 and CIC-DDoS 2017 [34]. UMNIDS includes flow-level and payload-based data with features that capture historical and temporal patterns in network traffic. This dataset includes around $170,000$ observations, $88$ features and 11 target variable, which implies if a record is normal or if it is an attack in the network along with the type of the attack. Some that are included are Worms, Bots, DDoS, Infiltration and SQL-Injection. The target variable includes a total of $44$ unique values, $43$ of which specify the kind of attack and the one remaining the non-malicious records. One of the main issues with this dataset is the lack of records for some attacks, like Heartbleed and Infiltrations.
@@ II-B Data Pre-Processing
The pre-processing procedure started by omitting the duplicate values, then encoding the label variable, testing different methods like binning for certain numeric variables with a small number of unique values. We performed an in-depth statistical analysis that included the discovering of linear and non-linear relationships, hypothesis testing for existence of normality through $Q$-Plots and goodness of fit tests like Kolmogorov-Smirnov and Anderson Darling, as well as regression analysis for some numeric features. We also examined the existence of outliers with IQR, Z-score, DBSCAN and Local-Outlier Factor and due the great number of outliers we preferred to keep them and use as a scaling method the Robust scaler, after we performed an $80/20$ train/test stratified split, with respect to the attack type variable.
@@ II-C Attacks Classification
For the classification of the tabular data, we selected baseline and top-tier models from different families. The two baseline models that we used for the multi-class classification task are Logistic Regression and the stochastic gradient descent-based SVM. The tree-based model that we utilized is a Decision Tree. Finally, the ensemble models are XGBoost, XGBoost with Random Forest, Random Forest, Extra Trees, Gradient Boosting, and Histogram-Based Gradient Boosting. Several hyperparameter optimization techniques were tested, such as grid-search, halving-grid-search, and Bayesian optimization, along with stratified $10$-fold cross validation. After the selection of the hyperparameters, each of the pre-mentioned models has been trained with stratified $10$-fold cross validation to observe their behavior without probability calibration. After that, we performed stratified $10$-fold cross validation with probability calibration, while holding a $4\%$ value range threshold for considering a model stable per metric, and obtained the results that follow for the training and testing set.
The metrics with which we evaluated the performance of each algorithm are F1, Precision, Recall, and Roc-Auc. From the results in Tab. I and Tab. II we observe the high performance of ensemble and tree-based models, and more specifically of XGBoost, XGBoost-RF, Random Forest, Extra Trees, and Decision Tree, while the baseline models, SVM and Logistic Regression fail. To highlight the impressive performance of the non-calibrated XGBoost, in terms of both stability and efficiency, we present the cross validation results for precision and recall in Fig. 1.
@@ III-B Training Adversarial Models
For the training of the adversarial networks we tested a few neural network architectures, we consider a total number of 2020 epochs for each of the above-mentioned algorithms. From the minimax procedure, for illustration purposes we can see below in Fig. 2 the discriminator’s and generator’s loss values for a few batches, for the Vanilla-GAN and WGAN, from which we observe the failure of Vanilla-GAN as the discriminator’s loss converges to zero and the stable training WGAN, along with indications of convergence:

### 27 Apr 2026 v3

@@ II-A Data Collection
For all the results that will be presented in the next sections we used the under-sampled version of the UMNIDS dataset, a newly created dataset, which was created by processing raw pcap files from CIC-IDS 2017, CIC-IOT 2023, UNSW-NB15 and CIC-DDoS 2017 [34]. UMNIDS includes flow-level and payload-based data with features that capture historical and temporal patterns in network traffic. This dataset includes around $170,000$ observations, $88$ features and $1$ target variable, which implies if a record is normal or if it is an attack in the network along with the type of the attack. Some that are included are Worms, Bots, DDoS, Infiltration and SQL-Injection. The target variable includes a total of $44$ unique values, $43$ of which specify the kind of attack and the one remaining the non-malicious records. One of the main issues with this dataset is the lack of records for some attacks, like Heartbleed and Infiltrations.
@@ II-B Data Pre-Processing
The pre-processing procedure started by omitting the duplicate values, then encoding the label variable, testing different methods like binning for certain numeric variables with a small number of unique values. We performed an in-depth statistical analysis that included the discovering of linear and non-linear relationships, hypothesis testing for existence of normality through QQ-Plots and goodness of fit tests like Kolmogorov-Smirnov and Anderson Darling, as well as regression analysis for some numeric features. We also examined the existence of outliers with IQR, Z-score, DBSCAN and Local-Outlier Factor and due the great number of outliers we preferred to keep them and use as a scaling method the Robust scaler, after we performed an $80/20$ train/test stratified split, with respect to the attack type variable.
@@ II-C Attacks Classification
For the classification of the tabular data, we selected baseline and top-tier models from different families. The two baseline models that we used for the multi-class classification task are Logistic Regression and the stochastic gradient descent-based SVM. The tree-based model that we utilized is a Decision Tree. The ensemble models are LightGBM, XGBoost, XGBoost with Random Forest, Random Forest, Extra Trees, Gradient Boosting, and Histogram-Based Gradient Boosting. Finally we also chose to use the AutoGluon, an ensemble of neural networks, extra trees, XGBoost, LightGBM and linear models. Several hyperparameter optimization techniques were tested, such as grid-search, halving-grid-search, and Bayesian optimization, along with stratified $10$-fold cross validation. After the selection of the hyperparameters, each of the pre-mentioned models has been trained with stratified $10$-fold cross validation to observe their behavior without probability calibration. After that, we performed stratified $10$-fold cross validation with probability calibration, while holding a $4\%$ value range threshold for considering a model stable per metric, and obtained the results that follow for the training and testing set.
The metrics with which we evaluated the performance of each algorithm are F1, Precision, Recall, and Roc-Auc. From the results in Tab. I and Tab. II we observe the high performance of ensemble and tree-based models, and more specifically of XGBoost, LightGBM, AutoGluon, XGBoost-RF, Random Forest, Extra Trees, and Decision Tree, while the baseline models, SVM and Logistic Regression fail. To highlight the impressive performance of the non-calibrated XGBoost, in terms of both stability and efficiency, we present the cross validation results for precision and recall in Fig. 1.
@@ III-B Training Adversarial Models
For the training of the adversarial networks we tested a few neural network architectures, we consider a total number of $20$ epochs for each of the above-mentioned algorithms. From the minimax procedure, for illustration purposes we can see below in Fig. 2 the discriminator’s and generator’s loss values for a few batches, for the Vanilla-GAN and WGAN, from which we observe the failure of Vanilla-GAN as the discriminator’s loss converges to zero and the stable training WGAN, along with indications of convergence:
@@ IV Conclusions
V Acknowledgments
This work has been supported by the University of Piraeus Research Center (UPRC).